# Why Johnny Signs with Sigstore: Examining Tooling as a Factor in Software Signing Adoption in the Sigstore Ecosystem

### v3

@@ 5.1.2 Sigstore’s Weaknessses
Enterprise Adoption Limitations 7
Rate Limiting Problems – T P3, P7, P10, P11
Lack of dedicated Support & Maintenance – T & P P11, P2
@@
Transparency Log Issues 6
Not Suitable for private Setup – T P2, P3, P6, P10, P13
Use in Air Gap Conditions – T P2, P3, P8
@@
Private Sigstore Instance Setup 5
Documentation – P P8, P9, P6
Limited Community Support – M P11
@@
Other Documentations and Usage Information IssuesP 3 — P1, P10, P13
Integration to Other Systems 3
@@
Gitlab & Jenkins – T P8
Other Unsupported technologies – T P12
@@
Enterprise Adoption Limitations: Many of our participants mentioned that Sigstore is not suited for large enterprise applications. The issues reported by participants include rate limiting—where Sigstore restricts the number of signatures that can be created per unit time—along with latency concerns, where the service’s turnaround time for a large volume of signatures is problematic. Additionally, participants highlighted the lack of dedicated support and maintenance teams due to Sigstore ’s open-source nature, as well as its unsuitability for organizations in regulated sectors for the same reason. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P8 (SAAS, SSC security): “I think the major challenge was that we were relying on the public Sigstore instance, and I don’t think it could meet our signing needs in terms of just the capacity of signatures we needed as an enterprise. We would constantly be getting rate limited by the service.” {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P3 (SAAS, SSC security tool): “For customers that are very large in scale, the biggest of the Fortune 100 or customers operating in very highly regulated environments, I think operating your own Sigstore instance within that air gap in a private environment could be valuable. And I think it depends on the amount of software you’re producing and the frequency with which you’re producing it.” The issues here were a mixture of all usability factors.
Transparency Log Issues: While the transparency log bundled with Sigstore was highlighted as a major advantage by participants, it was also a significant concern. Their concerns stemmed from the public nature of the log, where sensitive artifact metadata from company assets could be exposed in a publicly accessible record. Another common issue raised was the log’s usability in air-gapped environments—systems that are physically isolated from unsecured networks for security reasons. Additionally, participants noted the manual effort required to continuously monitor and manage the log. Most of these participants were in the process of experimenting with setting up the transparency log at the time of the interviews. Transparency log issues were mostly technological.
@@
[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P6 (security tooling, digital technology): “We’ve got some teams piloting Rekor, for instance, for different traceability usages and so on…I’m all for the transparency log, but we just have to be careful who it’s transparent to at what point in time.”
@@
Setting up Private Sigstore Instance : Sigstore’s customizability was a feature frequently utilized and discussed by participants; however, it was also associated with certain challenges. The primary issues identified included limited documentation on setup, scarce community usage information, and the infrastructure requirements for private deployments. Issues here were primarily related to participants’ need for support resources, which where macroenvironmental and social in nature. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P11 (internal security tooling, aerospace security): “There’s not a big enough community of practitioners, or help to help people deploy Fulcio, like on prem, or in their own infrastructure.”
@@
Other Documentation Issues: Overall, the documentation for Sigstore, along with practical usage information and support, was found to be insufficient. Participants noted that updates to the documentation lagged behind changes in the Sigstore product and that there was a lack of clarity regarding the use of different Sigstore components. P13 (security tooling, internet service) highlights this, “ I would say the documentation lacks a bit. Just I think if you look at the documentation, it was updated quite a long time ago, like the ReadMe file. And so if we could have these different options saying that, okay, if you are starting over, this is how you can do rekor and Fulcio.”
@@
Integration to Other Systems: Participants also expressed a desire for a wider range of compatible technologies beyond what Sigstore currently supports. Commonly mentioned integrations included other CI/CD platforms such as Jenkins and GitLab, as well as attestation storage databases. These limitations were often attributed to Sigstore being an emerging technology, as highlighted by P12 (SAAS, SSC security tool), “The weakness is not everything supports it. There’s a lot of weaknesses just around not just it’s being a newer technology”
@@
Fulcio Issues: Some Sigstore issues were also tied to its certificate authority’s (Fulcio) time stamping capabilities, and implementation of OIDC for keyless signing. Fulcio acts as an intermediary using OIDC identities to bind to a short-lived certificate, this was criticized by P3. P3 also commented on the initial lack of time stamping abilities for the initial versions of Fulcio. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P3 (SAAS, SSC security tool): “I think we do have some concerns about the way Fulcio operates as its own certificate authority. So we’ve been looking at things like OpenPubkey as something that removes that intermediary and would allow you to do identity-based signing directly against the OIDC provider. And then I think the other big thing that we did, because Sigstore originally didn’t support it, was time-stamping. Because we’re using Fulcio, we want to ensure that the signature was created when the certificate was valid. And so we use external time-stamping authority in order to be able to do that.”
@@
5.1.3 Effect of Identified Problems on Sigstore components used by Subjects
Table 5: Sigstore Functionalities Used by Subjects
Tool # of Subjects
Cosign (Signing & Verification CLI) 9
Certificate Authority (Fulcio) 9
Keyless Signing & Identity Manager (OIDC) 9
Transparency Log (Rekor) 8
Gitsign (Commit Signing) 4
Component Customization & Local Sigstore Deployment 2
We assessed the effect of participants’ reported problems with Sigstore on their usage of each component. Table 5 summarizes our subjects’ reported usage of Sigstore components and functionalities.
Key components like Cosign, Fulcio, and keyless signing saw high adoption, while local Sigstore deployment had low usage due to numerous reported issues. Several participants cited difficulties with documentation and setup, and as a result, only two successfully deployed private instances, likely due to these challenges (Table 4, section 5.1.2).
@@ 5.1.3 Effect of Identified Problems on Sigstore components used by Subjects
The Rekor transparency log presents a mixed case, with both high adoption and significant reported issues. Of the eight participants who used it, three were still in the pilot phase. Although Rekor was recognized as a key factor driving adoption due to its strengths, it also faced a high number of reported issues, indicating that, despite its benefits, challenges persist for some participants.
@@
Other Components Without Mention Though the participants in the survey shared experiences with user-facing components, discussion about other crucial elements was notably absent. In particular, there was very little mention of log witnessing [28], or monitoring [22] for either Fulcio or Rekor. This is remarkable, as transparency solutions require the existence of these parties to ensure log integrity and identify log misbehavior. We believe that targeted future work should identify possible gaps in user perception, as well as identify user stories so users verify log witnesses and monitor information.
5.2 RQ2: How and Why Practitioners Change Software Signing Tools
Table 6: Reasons Practitioners Choose or Switch to Sigstore Before Adoption.
Topics & Associated Examples Subjects
Human/Social Factors
Practitioners Contribute to Sigstore 6 – P2, P4, P6, P7, P8, P10
GPG Issues 6
Low adoption rates P1, P10
Key management issues & Other usability concerns P1, P6, P9, P12, P13
Steep learning curve P9, P12
Compatibility with newer technology P12
Notary Issues 2
Non-demand from customers P9
Compatibility with other tools P9
Lack of regular updates P9
Key & Identity Management P5
Proprietary Tool Issues 1
Difficult to setup P11
Technological Factors
Available Sigstore Functionalities 3 – P1, P5, P10
A transparency log, etc P5, P10
Compatibility to other Tools P1
Macroenvironmental Factors
Regulation & Standards 4 – P5, P6, P11, P13
Large User Community 1 – P8
Inherent Trust of Creators 1
Trust of CNCF products P3
Figure 3: Practitioners’ Evolution of Software Signing Tools. Most reported their organizations moved from no former tools, a form of GPG signing or notary to Sigstore. Most participants depicted with the flow from none to a sigstore implementation report that; they either initially used no tools, or have always used Sigstore.
We also tackle the usability of software signing tool especially as it relates to decisions that prompts the switch of practitioners in using different tools. In the context of our study, we asked participants what factors prior to adopting Sigstore influence practitioner’s to switch to Sigstore. The factors here are mostly Macroenvironmental (relating to regulations, user communities), Human (Previous experiences), and Technological (available capabilities of sigstore)
First, we present a Sankey flow diagram ( Figure 3) illustrating the evolution of signing tool usage among our participants’ teams up to the present. A significant portion of participants (6), as depicted transitioned from using no tools to adopting Sigstore —either they had always used Sigstore since the inception of their organization or continued using their initial signing tools alongside Sigstore. The most notable transition was from GPG-based signing implementations (3) to Sigstore, followed by a shift from Notary (2). Other reported transitions included Skopeo (1), unspecified internal tools (1), and a proprietary tool (1).
@@ 5.2 RQ2: How and Why Practitioners Change Software Signing Tools
The reasons most of our participants chose to switch to Sigstore stemmed from a combination of two key factors: issues related to a subpar user experience with previous tools and advantages associated with Sigstore. We summarize these factors in Table 6.
@@
Contributions to Sigstore: Being part of the Sigstore community as contributors was a major reason for adopting Sigstore. Some participants (e.g., P6) stated that their organizations have a policy of contributing to open-source communities they intend to adopt. Others noted that their organizations were relatively new compared to Sigstore—meaning that Sigstore had already existed before these organizations were established, and some of their members had contributed to Sigstore before joining their respective organizations. Additionally, some participants mentioned that key members of their organizations were directly involved in the initial development of the Sigstore project. P10 (POC software & security tool, cloud security) reflects this, “I’d say many of the early staff, the first 10 to 20 staff were involved in the Sigstore community. There’s not just our interest as a company in a Sigstore, but there’s a lot of personal ties to Sigstore.”
@@
Available Sigstore Functionalities: Sigstore’s advertised functionalities also influenced its adoption. While several participants were drawn to the transparency log (as highlighted in section 5.1.1), its integrability with existing technologies they were using at the time also played a role in their decision. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P5 (internal security tool, cloud security): “… Sigstore has other features, like you can use your OIDC identities, like your GitHub identities as well, to do the signing, as well as they also maintain a transparency log as well …” {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P1 (POC software, digital technology): “Factors I considered were, …, the existence of tools to use it.”
@@
[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P5 (internal security tool, cloud security): “…standards actually really helped to convince everyone that we should be doing this … But once a standard is put in place, once the requirements are set, then it just sped up the process of adopting these new technologies like Sigstore …”
User Community & Trust Of Sigstore Creators: The large user community of Sigstore was also a key reason for its adoption among participants. This aligns with feedback from some participants (P1, P10) who cited GPG’s smaller user base as a drawback. Additionally, some participants expressed inherent trust in the CNCF (Cloud Native Computing Foundation) security tooling community, which they had grown to rely on over time. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P3 (SAAS, SSC security tool): “We do defer often at the higher level projects to different foundations. So we would be more likely to trust something from the CNCF, the Cloud Native Computing Foundation or Sigstore rather than small little independent developer projects or things like that.”
@@
Notary-related concerns primarily revolved around customer demands, compatibility with other technologies, and infrequent updates. Key and identity management challenges were common to both GPG and Notary. Meanwhile, users of proprietary signing tools faced difficulties with complex setup processes.
6 Discussion
In this section, we highlight two reflections from our data.
The role of community in software signing tool adoption: A substantial number of participants reported that they adopted Sigstore because of their involvement as contributors to the Sigstore tooling community (section 5.2). This reflects a key macroenvironmental factor from our usability frameworks that influenced users’ decisions to adopt these tools. Similarly, several reported challenges were linked to the lack of community support in setting up private Sigstore instances and accessing usage information (section 5.1.2).
These findings highlight the critical role of community, as emphasized in prior research [72]. New adopters often depend on communities they trust, contribute to, or receive support from, which significantly shapes their decision to adopt such tools. Toolmakers should prioritize fostering and maintaining a strong community around their tools, as this can greatly influence adoption and success.

### v4

@@ V-A2 Sigstore’s Weaknessses
Enterprise Adoption Limitations 7 subjects
Rate Limiting Problems – T P3, P7, P14, P15
Lack of dedicated Support & Maintenance – T & P P15, P2
@@
Transparency Log Issues 6 subjects
Not Suitable for private Setup – T P2, P3, P6, P14, P17
Use in Air Gap Conditions – T P2, P3, P9
@@
Private Sigstore Instance Setup 5 subjects
Documentation – P & T P9, P12, P6
Limited Community Support – M P15
@@
Other Documentations and Usage Information IssuesP & T 3 — P1, P14, P17
Integration to Other Systems 3 subjects
@@
Gitlab & Jenkins – T P9
Other Unsupported technologies – T P16
@@
Macroenvironmental & Organizational
Enterprise Adoption Limitations: Many of our participants mentioned that Sigstore is not suited for large enterprise applications. The issues reported by participants include rate limiting—where Sigstore restricts the number of signatures that can be created per unit time—along with latency concerns, where the service’s turnaround time for a large volume of signatures is problematic. Additionally, participants highlighted the lack of dedicated support and maintenance teams due to Sigstore ’s open-source nature, as well as its unsuitability for organizations in regulated sectors for the same reason. {quoting} P7 (security tools, cloud security): “We were relying on the public Sigstore instance, and I don’t think it could meet our signing needs in terms of just the capacity of signatures we needed as an enterprise [rate limit].” {quoting} P3 (SAAS, SSC security tool): “For customers that are very large in scale, the biggest of the Fortune 100 or customers operating in very highly regulated environments, I think operating your own Sigstore instance within that air gap in a private environment could be valuable. And I think it depends on the amount of software you’re producing and the frequency with which you’re producing it.” The issues here were a mixture of all usability factors.
Technological
Transparency Log Issues: While the transparency log bundled with Sigstore was highlighted as a major advantage by participants, it was also a significant concern. Their concerns stemmed from the public nature of the log, where sensitive artifact metadata from company assets could be exposed in a publicly accessible record. Another common issue raised was the log’s usability in air-gapped [85] environments. Additionally, participants noted the manual effort required to continuously monitor and manage the log. Most of these participants were in the process of experimenting with setting up the transparency log at the time of the interviews.
@@
P6 (security tooling, digital technology): “We’ve got some teams piloting Rekor, for instance, for different traceability usages…I’m all for the transparency log, but we…have to be careful who it’s transparent to at what [time].”
@@
Setting up Private Sigstore Instance: Sigstore’s customizability was a feature frequently utilized by participants. However, it was also associated with certain challenges. The primary issues identified included limited documentation on setup, scarce community usage information, and the infrastructure requirements for private deployments. Issues here were primarily related to participants’ need for support resources, which were macroenvironmental and social in nature. {quoting} P15 (internal security tooling, aerospace security): “There’s not a big enough community of practitioners, or help to…deploy Fulcio, like on prem, or in their own infrastructure.”
@@
Other Documentation Issues: Overall, the documentation for Sigstore, along with practical usage information and support, was found to be insufficient. Participants noted that updates to the documentation lagged behind changes in the Sigstore product and that there was a lack of clarity regarding the use of different Sigstore components. P17 (security tooling, internet service) highlights this, “I would say the documentation lacks a bit. Just I think if you look at the documentation, it was updated quite a long time ago, like the ReadMe file. And so if we could have these different options saying that, okay, if you are starting over, this is how you can do rekor and Fulcio.”
@@
Integration to Other Systems: Participants also expressed a desire for a wider range of compatible technologies beyond what Sigstore currently supports. Commonly mentioned integrations included other CI/CD platforms such as Jenkins and GitLab, as well as attestation storage databases. These limitations were often attributed to Sigstore being an emerging technology, as highlighted by P16 (SAAS, SSC security tool), “The weakness is not everything supports it. There’s a lot of weaknesses around…its being a newer technology”
@@
Fulcio Issues: Some Sigstore issues were also tied to its certificate authority’s (Fulcio) time stamping capabilities, and implementation of OIDC for keyless signing. Fulcio acts as an intermediary using OIDC identities to bind to a short-lived certificate, this was criticized by P3. P3 also commented on the lack of timestamping support in early versions of Fulcio. {quoting} P3 (SAAS, SSC security tool): “We do have some concerns about the way Fulcio operates as its own certificate authority. So we’ve been looking at things like OpenPubkey as something that removes that intermediary and would allow you to do identity-based signing directly against the OIDC provider. And then I think the other big thing that we did, because Sigstore originally didn’t support it, was time-stamping. We want to ensure that the signature was created when the certificate was valid.”
@@
V-A3 Effect of Identified Problems on Sigstore components used by Subjects
We assessed the effect of participants’ reported problems with Sigstore on their usage of each component. By component, nine subjects use Cosign, nine use the Fulcio certificate authority, and nine use the OIDC keyless signing and identity manager. Eight use the Rekor transparency log. Only four use Gitsign and only two make use of customized components and local Sigstore deployments.
@@ V-A3 Effect of Identified Problems on Sigstore components used by Subjects
Among these, our data shed the most light on local deployments and the Rekor log. For local deployment, several participants cited difficulties with documentation and setup, and as a result, only two successfully deployed private instances, likely due to these challenges (Table IV, section V-A2). The Rekor transparency log presents a mixed case, with both high adoption and significant reported issues. Of the eight participants who used it, three were still in the pilot phase. Although Rekor was recognized as a key factor driving adoption due to its strengths, it also faced a high number of reported issues, indicating that, despite its benefits, challenges persist for some participants.
@@
Other Components Without Mention Though the participants in the survey shared experiences with user-facing components, discussion about other elements was notably absent. In particular, there was very little mention of log witnessing [86], or monitoring [87] for either Fulcio or Rekor. This is remarkable, as transparency solutions require the existence of these parties to ensure log integrity and identify log misbehavior. We believe that targeted future work should identify possible gaps in user perception, as well as identify user stories so users verify log witnesses and monitor information.
V-B RQ2: How & Why Practitioners Change Tools
V-B1 Drivers of Adoption Among Current Sigstore Users
For participants who adopted Sigstore, we asked which factors—prior to adoption—influenced their decision (or their organization’s decision) to switch. Their responses fell into three main categories: macroenvironmental factors (e.g., regulations and user communities), human/social factors (e.g., prior experience with signing tools), and technological factors (e.g., Sigstore’s unique capabilities).
@@ V-B1 Drivers of Adoption Among Current Sigstore Users
Most of our participants chose to switch to Sigstore due to poor experiences with previous tools and due to perceived advantages associated with Sigstore. We summarize these factors in Table X.
@@
Contributions to Sigstore: Being part of the Sigstore community as contributors was a major reason for adopting Sigstore. Some participants (e.g., P6) stated that their organizations have a policy of contributing to open-source communities they intend to adopt. Others noted that their organizations were relatively new compared to Sigstore—meaning that Sigstore had already existed before these organizations were established, and some of their members had contributed to Sigstore before joining their respective organizations. Additionally, some participants mentioned that key members of their organizations were directly involved in the initial development of the Sigstore project. P14 (POC software & security tool, cloud security) reflects this, “I’d say many of the early staff, the first 10 to 20 staff were involved in the Sigstore community. There’s not just our interest as a company in a Sigstore, but there’s a lot of personal ties to Sigstore.”
@@
Available Sigstore Functionalities: Sigstore’s functionalities also influenced its adoption. While several participants were drawn to the transparency log (as highlighted in section V-A1). {quoting} P5 (internal security tool, cloud security): “Sigstore has other features, like you can use your OIDC identities, like your GitHub identities as well, to do the signing, [and] they also maintain a transparency log.”
Integration with existing technologies also affected their decision. As P1 (POC software, digital technology) said, “Factors I considered were, …, the existence of tools to use it.”
@@
P5 (internal security tool, cloud security): “…standards actually really helped to convince everyone that we should be doing this … But once a standard is put in place, once the requirements are set, then it just sped up the process.”
User Community & Trust Of Sigstore Creators: The large user community of Sigstore also motivated its adoption among participants. Conversely, participants P1 and P14 cited GPG’s smaller user base as a drawback. Additionally, some participants expressed trust in industry consortia such as CNCF (Cloud Native Computing Foundation) {quoting} P3 (SAAS, SSC security tool): “We do defer often at the higher-level projects to different foundations. So we would be more likely to trust something from the CNCF…[than] independent developer projects.”
@@
Notary-related issues primarily concerned customer demands, compatibility with other technologies, and infrequent updates. Key and identity management challenges were common to both GPG and Notary. Meanwhile, users of proprietary signing tools faced difficulties with complex setup processes.
V-B2 Barriers and Motivators Spurring Consideration of Sigstore Among Non-Sigstore Users
TABLE V: Non Sigstore Users: Reasons for Not Adopting Sigstore. This table summarizes the reasons non-users gave for not adopting Sigstore, grouped by their current signing tool and mapped to our usability framework theme.
Tool Subjects Reason (Topic)
Internal tool P8, P11 (1) Third-party security risk from tool T – P11, P8
(2) Organizational Use cases O – S8
(3) Centralizing Organization’s code signing O – P8
(4) Organization’s Open Source Policy O – P8
Notary-V1 P10 (1) Tool owned by Organization O – S10
PGP P13 (1) Sigstore Privacy concerns T – P13
(2) Trust in Sigstore’s Maintainers M – P13
In this section we highlight reasons mentioned by practitioners that have stopped them (or their organization) from switching to Next-gen tools like sigstore. We also report issues experienced by these non sigstore users that have prompted a consideration (or in some cases partial adoption of these next-gen tools).
Barriers to Considering Sigstore: As highlighted in Table V, most reasons for non-adoption were organizational. {quoting} P8 (internal security tool & cloud APIs, internet services): “When it comes to open source projects and open source stuff, our Organization is very skeptical. Policy and everything drives very much teams away from that sort of stuff…There’s also the ongoing risk of having a product that’s built outside continually shipping code and being integrated into this product, where it deals with sensitive material. So we do evaluate these things, but the scrutiny against code written outside of the company is extreme, and with good reason given the context of all the supply chain stuff.” Other notable issues arose from technical concerns, including privacy risks associated with Sigstore’s public instance and the inherent risk of relying on a third-party tool maintained by open-source contributors.
Factors Motivating Consideration for sigstore: When practitioners were asked about issues with their current tools and potential fixes, most pointed to existing tooling problems as the primary driver for considering alternatives—such as incorporating next-gen features into internal tools or partially adopting new solutions. Although some subjects reported drawing design inspiration from next-generation workflows, none specifically mentioned Sigstore. P11 (firmware & testing software, social technology) states, “. I’m sure at some point in time it was influenced by an open source project before they adopted it and made the right tweaks internally.” Lack of transparency and integration of these tools to other platforms were also common. We summarize these factors in Table XI.
VI Discussion
VI-A Lessons for Next-Generation Tools
Next-generation signing tools offer unparalleled flexibility by integrating with external platforms, and services. However, our participants consistently highlighted integration challenges as a major obstacle.
VI-A1 Automated Cross-Platform Integrations
To fully leverage the flexibility of next-gen tools, practitioners need integration modules for popular CI/CD systems, registries, and orchestration platforms. Missing integration and sparse documentation forces users to implement custom connectors, increasing the risk of misconfiguration and potential security vulnerabilities that undermine the very guarantees signing is meant to provide.
VI-A2 Need for Inter-Platform Standards
Standardization of information exchange is a standard problem for new classes of software infrastructure e.g., , SBOM [88]. These problems often harm integration capabilities. Next-generation signing tools would gain from common exchange standards—for example, uniform schemas for transparency logs or identity assertions. Projects like Diverfiy [89], which harmonize identity provisioning across multiple OIDC providers, are a promising start. Similar efforts are needed to streamline data flows between signing components, transparency services, and external tooling ecosystems.
VI-A3 Verification Workflow
A key security property in a software supply chain is the transparency of actors and artifacts, which requires verifying signer identities [4]. Previous studies have highlighted challenges in signature verification across both organizational and open-source contexts [39, 5]. Next-generation tools like Sigstore integrate multiple verification methods, such as the Cosign CLI and the transparency log, but most participants still reported usability issues with the transparency log, and documentation indicates a lack of automation for other verification methods. Further automating the verification workflow would enhance the security benefits of software signing with this class of tools.
VI-B Usability Patterns of Sigstore vs Non-Sigstore users
Our results reveal different drives in adoption motivations and shared usability concerns between Sigstore users and non-users. Sigstore adopters cited contributions to the Sigstore project (Human/Social factor) and macroenvironmental factors—alongside negative experiences with legacy tools—as their primary drivers. In contrast, non-users emphasized organizational considerations when choosing their signing solution.
Despite these differences, both groups reported similar technological usability issues, particularly around integration with other systems and infrastructure transparency. Notably, non-users frequently highlighted signer identification challenges, whereas Sigstore users did not—underscoring the clear usability advantage of next-generation tooling.
VI-C Privacy Concerns